# Understanding Nonlinear Implicit Bias via Region Counts in Input Space

### v1

@@ Abstract
footnotetext: Equal contributionfootnotetext: Corresponding author
@@ 1 Introduction
One mystery in deep neural networks lies in their ability to generalize, despite having significantly more learnable parameters than the number of training examples (zhang2017understanding, ). The choice of network architectures, including factors such as nonlinearity, depth, and width, along with training procedures like initialization, optimization algorithms, and loss functions, can result in vastly diverse generalization performance (sutskever2013importance, ; smith2017don, ; wilson2017marginal, ; li2019towards, ). The varied generalization abilities exhibited by neural networks are often explained by many researchers through the theory of implicit bias (brutzkus2017sgd, ; soudry2018implicit, ). Implicit bias refers to the inherent tendencies in how neural networks learn and generalize from the training data, even without explicit regularizations or constraints.
The implicit bias of linear neural networks has been extensively studied. One of the classical setting is linear classification with logistic loss. brutzkus2017sgd ; soudry2018implicit ; arora2019implicit show that the parameter converges to the direction that maximizes the $L_{2}$ margin. For regression problems, it is proved that gradient descent or stochastic gradient descent converges to a parameter that is closest to the initialization in terms of $L_{2}$ norm (gunasekar2018characterizing, ). The results of the linear regression model can be extended to deep linear neural networks by generalizing the definition of min-norm and max-margin solutions (ji2018gradient, ; vaskevicius2019implicit, ; woodworth2020kernel, ).
Compared to linear models, defining implicit biases in non-linear networks poses significant challenges. One line of work studies homogeneous networks and demonstrates that gradient flow solutions converge to a KKT point of the max-margin problem (lyu2019gradient, ; ji2020directional, ; wang2021implicit, ; jacot2022feature, ). Further research extends this analysis, showing that gradient flow converges to a max-margin solution under various norms (ongie2019function, ; chizat2020implicit, ). Other studies focus on describing the implicit bias of neural networks using sharpness, such as (foret2020sharpness, ; montufar2022sharp, ; andriushchenko2023modern, ).
@@
Motivated by the above studies, we focus on leveraging decision boundaries in the input space to characterize implicit bias. We adopt a metric called region count, which measures the average number of connected components in the decision regions of a predictor (see Figure 1). This idea was initially introduced by somepalli2022can , who used it under the name fragmentation score to study the double descent phenomenon by counting decision regions on two-dimensional planes defined by training examples. In contrast, we extend this idea beyond 2D subspaces and generalize the region count to arbitrary-dimensional input subspaces, which enables a more comprehensive geometric characterization.
To compute region count, we project the input space onto low-dimensional subspaces, use the trained network to predict the labels of points within these subspaces, and count the number of connected regions that share the same predicted label. Unlike previous studies on linear region count (hanin2019complexity, ; hanin2019deep, ; safran2022effective, ), which focus on activation patterns and representational capacity, the metric of region count is label-dependent and reflects the functional behavior of the classifier. We find that region count correlates strongly with the generalization gap—defined as the difference between training and test errors. As illustrated in Figure 1, models with fewer decision regions tend to generalize better.
@@ 2 Related Works
The implicit bias in linear neural networks are thoroughly investigated in recent works. For linear logistic regression on linearly separable data, full-batch gradient descent converges in the direction of the maximum margin solution (soudry2018implicit, ). This foundational work has various follow-ups, including extensions to non-linearly-separable data (ji2018risk, ; ji2019implicit, ), stochastic gradient descent (nacson2019stochastic, ), and other loss functions and optimizers (gunasekar2018characterizing, ).
These findings in linear logistic regression are generalized to deep linear networks. For fully-connected neural networks with linear separable data, ji2018gradient shows that the direction of weight also converges to $L_{2}$ max-margin solution. For linear diagonal networks, the gradient flow maximizes the margin with respect to a specific quasi-norm that is related to the depth of network (gunasekar2018implicit, ; woodworth2020kernel, ; pesme2021implicit, ), leading to a bias towards sparse linear predictors as the depth goes to infinity. This sparsity bias also exists in linear convolutional networks (gunasekar2018implicit, ; yun2020unifying, ).
@@
The non-linearity of modern non-linear neural networks pose challenges to studying its implicit bias. Initial works in this area (lyu2019gradient, ; ji2020directional, ) focus on homogeneous networks. These studies show that with exponentially-tailed classification losses, both gradient flow and gradient descent converge directionally to a KKT point in a maximum-margin problem. wang2021implicit consider a more general setup that includes different optimizers and prove that both Adam and RMSProp are capable of maximizing the margin in neural networks while Adagrad is not. ongie2019function ; chizat2020implicit showcased a bias towards maximizing the margin in a variation norm for infinite-width two-layer homogeneous networks. lyu2021gradient ; jacot2022feature identified margin maximization in two-layer Leaky-ReLU networks trained with linearly separable and symmetric data. More recent investigations into non-linear neural networks, such as (jacot2022implicit, ), focus on the homogeneity of the non-linear layer, demonstrating an implicit bias characterized by a novel non-linear rank.
@@
Prior work has extensively studied the number of linear regions in ReLU networks (hanin2019complexity, ; hanin2019deep, ), where a linear region refers to a set of inputs sharing the same activation pattern. For example, safran2022effective showed that for a two-layer ReLU network with width $r$, gradient flow converges directionally to a network with at most $O(r)$ linear regions. Other studies (serra2018bounding, ; cai2023getting, ) demonstrated that increasing the number of linear regions can enhance representational capacity and potentially improve performance. However, linear regions are defined independently of output labels and thus primarily reflect the expressiveness of the network rather than its generalization behavior.
In contrast, decision regions are defined as connected subsets of the input space that correspond to the same predicted label. nguyen2018neural investigated decision regions over the full input space and showed that, under certain conditions, each class tends to form a single connected region. However, connectivity in the full space does not necessarily imply connectivity in a subspace—for instance, two points may be connected in 3D but disconnected in a 2D slice. Subspace analysis thus reveals a richer and more nuanced structure of decision boundaries.
The paper (somepalli2022can, ) is the most relevant to our work, which is the first to systematically analyze the number of decision regions in input subspaces. They introduce the fragmentation score, defined as the average number of decision regions over 2D planes determined by triplets of training points. Their study focuses on illustrating how fragmentation varies with network width and its connection to the double descent phenomenon. While their approach provides important empirical observations, the analysis remains largely qualitative and limited to specific architectural variations.
@@ 3 Motivation
$d(\cdot)$ and $\gamma(\cdot)$ are commonly used indicator for implicit bias of linear models (soudry2018implicit, ; ji2018risk, ). However, both of them are not invariant to network reparameterization. We can construct a different set of network weight parameters by scaling the parameters as $\hat{W}=\{2W_{1},\frac{1}{2}W_{2},\cdots,W_{l}\}$, such that $f_{W}=f_{\hat{W}}$ but $d(W)\neq d(\hat{W})$ in general. Similarly, we can scale the last layer weights and get $\tilde{W}=\{W_{1},\cdots,2W_{l}\}$, such that $\gamma(\tilde{W})\neq\gamma(W)$, but $\text{argmax}_{i}f_{W}(x)=\text{argmax}_{i}f_{\tilde{W}}(x)$. This reparameterization trick also works for more complicated norm-based and margin-based generalization metric in (jiang2019fantastic, ), or the sharpness-based metrics (andriushchenko2023modern, ).
We numerically investigate whether they are effective measures, by training a ResNet18 on CIFAR-10 dataset, using different hyperparameters as in Table 1. The results are presented in Figure 2, which indicates that these measures have a low correlation with the generalization gap in the deep learning regime. One could choose other definitions of norms to achieve stronger correlations, but such choices are often problem-specific and require domain expertise, as discussed in (jiang2019fantastic, ).
The definition of margin may also be improved to the input-space margin, i.e., the $\ell_{2}$ distance of input data $x$ to decision boundary defined by the classifier, which is able to characterize the quality and robustness of the classifier. This metric is invariant to reparameterization and therefore more intrinsic to the underlying classifier. However, due to the highly nonconvex loss landscape, the input-space margin is NP-complete to compute and even hard to approximate (katz2017reluplex, ; weng2018towards, ). Therefore, quantitatively analyzing the decision boundary of a neural network and characterizing its implicit bias remains a challenge.
@@ 4.1 Definition of Region Counts
Definition 2 (Maximally Connected Region).
@@
Definition 3 (Region Count).
@@ 4.2 Estimating Region Counts
Calculating the region count in the original high-dimensional input space can be computationally intractable. We follow the estimation method in (somepalli2022can, ), use a computationally efficient surrogate by calculating the region counts on low dimensional subspace spanned by training data points.
Definition 4 (Region Count in $d$-Dimensional Subspace).
@@ 5 Region Counts Correlate with Generalization Gaps
We conduct image classification experiments on the CIFAR-10 dataset, using different architectures, including ResNet18 (he2016deep, ), EfficientNetB0 (tan2019efficientnet, ), and SeNet18 (hu2018squeeze, ). Results on other architectures are deferred to ablation studies. We vary the hyperparameters for training, such as learning rate, batch size and weight decay coefficient, whose numbers are reported in Table 1. The region count is calculated using randomly generated 1D hyperplanes, as described in Example 2. We run each experiment 100 times and report the average number.
@@ 6.1 The Bias from Training Hyperparameters
Previous studies (keskar2016large, ; jastrzkebski2017three, ; hoffer2017train, ; novak2018sensitivity, ; lewkowycz2020large, ) find that certain hyperparameters, such as a large learning rate and a small batch size, can improve the generalization of the neural network. Our observations provide a possible explanation: these good hyperparameter choices lead to a reduced region count. Such simplicity bias can decrease the generalization gap of neural networks.
@@ 6.2 Theoretical Explanations
The validity of Assumption 1 comes from the fact that the bifurcation zone (bertoin2021numerical, ) of ReLU neural networks, which contains its non-differentiable points, has Lebesgue measure zero (bolte2020mathematical, ; bolte2021conservative, ; bianchi2022convergence, ). Therefore, if the distribution of weights are absolutely continuous with respect to the Lebesgue measure, the bifurcation zone can be avoided with probability one. We conjecture that it can be proved rigorously, but leave it as an assumption since the proof diverges from the main content in this paper.
The next assumption characterizes the sharpness along the training trajectory. This is actually from the well-known edge of stability phenomenon (cohen2020gradient, ; damian2022self, ; arora2022understanding, ; ahn2024learning, ), which states that the sharpness of neural networks, characterized by the $\ell_{2}$ norm of the Hessian matrix, hovers around $\frac{2}{\eta}$.
Assumption 2 (Edge of Stability).
@@
Theorem 1.
Under Assumption 1 and 2, we have that for neural net weights $W_{t}$ at training step $t\geq T$, with probability one, the average region count $R(X,X^{\prime},W_{t})$ for random training data point $X,X^{\prime}$ can be bounded as:
$\mathbb{E}_{X,X^{\prime}}[R(X,X^{\prime},W_{t})]=\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}R(x_{i},x_{j},W_{t})\leq O\left(\frac{N}{r^{2}\eta}\right).$
@@
Note that this lemma brings in an additional $N$ in the denominator, which leads to a $N$-dependent bound in Theorem 1. The $N$ dependency is actually tight in worst case analysis, by considering $N$ points on a line with alternating labels. We conjecture that the $N$-dependency can be optimized under further structural assumptions on the data distribution, and leave it for further investigations. Equipped with these two lemmas, Theorem 1 is a consequence of the sharpness condition in Assumption 2.
@@ 7 Ablation Studies
We first examine the influence of neural network architectures and datasets on our results. We provide additional results on various neural network architectures such as ResNet34 (he2016deep, ), VGG19 (simonyan2014very, ), MobileNetV2 (sandler2018mobilenetv2, ), ShuffleNetV2 (ma2018shufflenet, ), RegNet200MF (radosavovic2020designing, ), and SimpleDLA (yu2018deep, ). We also use various datasets such as CIFAR-100 (krizhevsky2009learning, ) and ImageNet (deng2009imagenet, ). Region counts and generalization gaps are evaluated across various learning rates, batch sizes, and weight decay parameters as listed in Table 1.
@@
Mixup (zhang2017mixup, ) is a data augmentation technique that creates training samples by linearly interpolating pairs of input data and their corresponding labels. We train a ResNet-18 model using mixup, with other hyperparameters in Table 1. The plot in Figure 6 illustrates that Mixup induces smoother decision boundaries with smaller region count and has a better generalization performance. Random crop and random horizontal flip is another way to enhance the diversity of the training dataset. We apply random crop of size 32$\times$32 with padding 4 and random horizontal flip with a probability of 0.5 as data augmentations. As depicted in Figure 7, we observe that compared with mixup, random crop and random flip result in a more evident vertical shift in the performance curve.

### v2

@@ Abstract
Machine Learning, ICML
@@ 1 Introduction
One mystery in deep neural networks lies in their ability to generalize, despite having significantly more learnable parameters than the number of training examples (Zhang et al., 2017a). The choice of network architectures, including factors such as nonlinearity, depth, and width, along with training procedures like initialization, optimization algorithms, and loss functions, can result in vastly diverse generalization performance (Sutskever et al., 2013; Smith et al., 2017; Wilson et al., 2017; Li et al., 2019). The varied generalization abilities exhibited by neural networks are often explained by many researchers through the theory of implicit bias (Brutzkus et al., 2017; Soudry et al., 2018). Implicit bias refers to the inherent tendencies in how neural networks learn and generalize from the training data, even without explicit regularizations or constraints.
The implicit bias of linear neural networks has been extensively studied. One of the classical setting is linear classification with logistic loss. Brutzkus et al. (2017); Soudry et al. (2018); Arora et al. (2019) show that the parameter converges to the direction that maximizes the $L_{2}$ margin. For regression problems, it is proved that gradient descent or stochastic gradient descent converges to a parameter that is closest to the initialization in terms of $L_{2}$ norm (Gunasekar et al., 2018a). The results of the linear regression model can be extended to deep linear neural networks by generalizing the definition of min-norm and max-margin solutions (Ji & Telgarsky, 2018a; Vaskevicius et al., 2019; Woodworth et al., 2020).
Compared to linear models, defining implicit biases in non-linear networks poses significant challenges. One line of work studies homogeneous networks and demonstrates that gradient flow solutions converge to a KKT point of the max-margin problem (Lyu & Li, 2019; Ji & Telgarsky, 2020; Wang et al., 2021; Jacot et al., 2022). Further research extends this analysis, showing that gradient flow converges to a max-margin solution under various norms (Ongie et al., 2019; Chizat & Bach, 2020). Other studies focus on describing the implicit bias of neural networks using sharpness, such as (Foret et al., 2020; Montúfar et al., 2022; Andriushchenko et al., 2023).
@@
Motivated by the above studies, we focus on leveraging decision boundaries in the input space to characterize implicit bias. We adopt a metric called region count, which measures the average number of connected components in the decision regions of a predictor (see Figure 1). This idea was initially introduced by Somepalli et al. (2022), who used it under the name fragmentation score to study the double descent phenomenon by counting decision regions on two-dimensional planes defined by training examples. In contrast, we extend this idea beyond 2D subspaces and generalize the region count to arbitrary-dimensional input subspaces, which enables a more comprehensive geometric characterization.
To compute region count, we project the input space onto low-dimensional subspaces, use the trained network to predict the labels of points within these subspaces, and count the number of connected regions that share the same predicted label. Unlike previous studies on linear region count (Hanin & Rolnick, 2019a, b; Safran et al., 2022), which focus on activation patterns and representational capacity, the metric of region count is label-dependent and reflects the functional behavior of the classifier. We find that region count correlates strongly with the generalization gap—defined as the difference between training and test errors. As illustrated in Figure 1, models with fewer decision regions tend to generalize better.
@@ 2 Related Works
The implicit bias in linear neural networks are thoroughly investigated in recent works. For linear logistic regression on linearly separable data, full-batch gradient descent converges in the direction of the maximum margin solution (Soudry et al., 2018). This foundational work has various follow-ups, including extensions to non-linearly-separable data (Ji & Telgarsky, 2018b, 2019), stochastic gradient descent (Nacson et al., 2019), and other loss functions and optimizers (Gunasekar et al., 2018a).
These findings in linear logistic regression are generalized to deep linear networks. For fully-connected neural networks with linear separable data, Ji & Telgarsky (2018a) shows that the direction of weight also converges to $L_{2}$ max-margin solution. For linear diagonal networks, the gradient flow maximizes the margin with respect to a specific quasi-norm that is related to the depth of network (Gunasekar et al., 2018b; Woodworth et al., 2020; Pesme et al., 2021), leading to a bias towards sparse linear predictors as the depth goes to infinity. This sparsity bias also exists in linear convolutional networks (Gunasekar et al., 2018b; Yun et al., 2020).
@@
The non-linearity of modern non-linear neural networks pose challenges to studying its implicit bias. Initial works in this area (Lyu & Li, 2019; Ji & Telgarsky, 2020) focus on homogeneous networks. These studies show that with exponentially-tailed classification losses, both gradient flow and gradient descent converge directionally to a KKT point in a maximum-margin problem. Wang et al. (2021) consider a more general setup that includes different optimizers and prove that both Adam and RMSProp are capable of maximizing the margin in neural networks while Adagrad is not. Ongie et al. (2019); Chizat & Bach (2020) showcased a bias towards maximizing the margin in a variation norm for infinite-width two-layer homogeneous networks. Lyu et al. (2021); Jacot et al. (2022) identified margin maximization in two-layer Leaky-ReLU networks trained with linearly separable and symmetric data. More recent investigations into non-linear neural networks, such as (Jacot, 2022), focus on the homogeneity of the non-linear layer, demonstrating an implicit bias characterized by a novel non-linear rank.
@@
Prior work has extensively studied the number of linear regions in ReLU networks (Hanin & Rolnick, 2019a, b), where a linear region refers to a set of inputs sharing the same activation pattern. For example, Safran et al. (2022) showed that for a two-layer ReLU network with width $r$, gradient flow converges directionally to a network with at most $O(r)$ linear regions. Other studies (Serra et al., 2018; Cai et al., 2023) demonstrated that increasing the number of linear regions can enhance representational capacity and potentially improve performance. However, linear regions are defined independently of output labels and thus primarily reflect the expressiveness of the network rather than its generalization behavior.
In contrast, decision regions are defined as connected subsets of the input space that correspond to the same predicted label. Nguyen et al. (2018) investigated decision regions over the full input space and showed that, under certain conditions, each class tends to form a single connected region. However, connectivity in the full space does not necessarily imply connectivity in a subspace—for instance, two points may be connected in 3D but disconnected in a 2D slice. Subspace analysis thus reveals a richer and more nuanced structure of decision boundaries.
The paper (Somepalli et al., 2022) is the most relevant to our work, which is the first to systematically analyze the number of decision regions in input subspaces. They introduce the fragmentation score, defined as the average number of decision regions over 2D planes determined by triplets of training points. Their study focuses on illustrating how fragmentation varies with network width and its connection to the double descent phenomenon. While their approach provides important empirical observations, the analysis remains largely qualitative and limited to specific architectural variations.
@@ 3 Motivation
$d(\cdot)$ and $\gamma(\cdot)$ are commonly used indicator for implicit bias of linear models (Soudry et al., 2018; Ji & Telgarsky, 2018b). However, both of them are not invariant to network reparameterization. We can construct a different set of network weight parameters by scaling the parameters as $\hat{W}=\{2W_{1},\frac{1}{2}W_{2},\cdots,W_{l}\}$, such that $f_{W}=f_{\hat{W}}$ but $d(W)\neq d(\hat{W})$ in general. Similarly, we can scale the last layer weights and get $\tilde{W}=\{W_{1},\cdots,2W_{l}\}$, such that $\gamma(\tilde{W})\neq\gamma(W)$, but $\text{argmax}_{i}f_{W}(x)=\text{argmax}_{i}f_{\tilde{W}}(x)$. This reparameterization trick also works for more complicated norm-based and margin-based generalization metric in (Jiang et al., 2019), or the sharpness-based metrics (Andriushchenko et al., 2023).
We numerically investigate whether they are effective measures, by training a ResNet18 on CIFAR-10 dataset, using different hyperparameters as in Table 1. The results are presented in Figure 2, which indicates that these measures have a low correlation with the generalization gap in the deep learning regime. One could choose other definitions of norms to achieve stronger correlations, but such choices are often problem-specific and require domain expertise, as discussed in (Jiang et al., 2019).
The definition of margin may also be improved to the input-space margin, i.e., the $\ell_{2}$ distance of input data $x$ to decision boundary defined by the classifier, which is able to characterize the quality and robustness of the classifier. This metric is invariant to reparameterization and therefore more intrinsic to the underlying classifier. However, due to the highly nonconvex loss landscape, the input-space margin is NP-complete to compute and even hard to approximate (Katz et al., 2017; Weng et al., 2018). Therefore, quantitatively analyzing the decision boundary of a neural network and characterizing its implicit bias remains a challenge.
@@ 4.1 Definition of Region Counts
Definition 4.2 (Maximally Connected Region).
@@
Definition 4.3 (Region Count).
@@ 4.2 Estimating Region Counts
Calculating the region count in the original high-dimensional input space can be computationally intractable. We follow the estimation method in (Somepalli et al., 2022), use a computationally efficient surrogate by calculating the region counts on low dimensional subspace spanned by training data points.
Definition 4.4 (Region Count in $d$-Dimensional Subspace).
@@ 5 Region Counts Correlate with Generalization Gaps
We conduct image classification experiments on the CIFAR-10 dataset, using different architectures, including ResNet18 (He et al., 2016), EfficientNetB0 (Tan & Le, 2019), and SeNet18 (Hu et al., 2018). Results on other architectures are deferred to ablation studies. We vary the hyperparameters for training, such as learning rate, batch size and weight decay coefficient, whose numbers are reported in Table 1. The region count is calculated using randomly generated 1D hyperplanes, as described in Example 2. We run each experiment 100 times and report the average number.
@@ 6.1 The Bias from Training Hyperparameters
Previous studies (Keskar et al., 2016; Jastrzkebski et al., 2017; Hoffer et al., 2017; Novak et al., 2018; Lewkowycz et al., 2020) find that certain hyperparameters, such as a large learning rate and a small batch size, can improve the generalization of the neural network. Our observations provide a possible explanation: these good hyperparameter choices lead to a reduced region count. Such simplicity bias can decrease the generalization gap of neural networks.
@@ 6.2 Theoretical Explanations
The validity of Assumption 6.1 comes from the fact that the bifurcation zone (Bertoin et al., 2021) of ReLU neural networks, which contains its non-differentiable points, has Lebesgue measure zero (Bolte & Pauwels, 2020, 2021; Bianchi et al., 2022). Therefore, if the distribution of weights are absolutely continuous with respect to the Lebesgue measure, the bifurcation zone can be avoided with probability one. We conjecture that it can be proved rigorously, but leave it as an assumption since the proof diverges from the main content in this paper.
The next assumption characterizes the sharpness along the training trajectory. This is actually from the well-known edge of stability phenomenon (Cohen et al., 2020; Damian et al., 2022; Arora et al., 2022; Ahn et al., 2024), which states that the sharpness of neural networks, characterized by the $\ell_{2}$ norm of the Hessian matrix, hovers around $\frac{2}{\eta}$.
Assumption 6.2 (Edge of Stability).
@@
Theorem 6.3.
Under Assumption 6.1 and 6.2, we have that for neural net weights $W_{t}$ at training step $t\geq T$, with probability one, the average region count $R(X,X^{\prime},W_{t})$ for random training data point $X,X^{\prime}$ can be bounded as:
$\mathbb{E}_{X,X^{\prime}}[R(X,X^{\prime},W_{t})]$
$=$ $\frac{1}{N^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}R(x_{i},x_{j},W_{t})\leq O\left(\frac{N}{r^{2}\eta}\right).$
@@
Note that this lemma brings in an additional $N$ in the denominator, which leads to a $N$-dependent bound in Theorem 6.3. The $N$ dependency is actually tight in worst case analysis, by considering $N$ points on a line with alternating labels. We conjecture that the $N$-dependency can be optimized under further structural assumptions on the data distribution, and leave it for further investigations. Equipped with these two lemmas, Theorem 6.3 is a consequence of the sharpness condition in Assumption 6.2.
@@ 7 Ablation Studies
We first examine the influence of neural network architectures and datasets on our results. We provide additional results on various neural network architectures such as ResNet34 (He et al., 2016), VGG19 (Simonyan & Zisserman, 2014), MobileNetV2 (Sandler et al., 2018), ShuffleNetV2 (Ma et al., 2018), RegNet200MF (Radosavovic et al., 2020), and SimpleDLA (Yu et al., 2018). We also use various datasets such as CIFAR-100 (Krizhevsky et al., 2009) and ImageNet (Deng et al., 2009). Region counts and generalization gaps are evaluated across various learning rates, batch sizes, and weight decay parameters as listed in Table 1.
@@
Mixup (Zhang et al., 2017b) is a data augmentation technique that creates training samples by linearly interpolating pairs of input data and their corresponding labels. We train a ResNet-18 model using mixup, with other hyperparameters in Table 1. The plot in Figure 6 illustrates that Mixup induces smoother decision boundaries with smaller region count and has a better generalization performance. Random crop and random horizontal flip is another way to enhance the diversity of the training dataset. We apply random crop of size 32$\times$32 with padding 4 and random horizontal flip with a probability of 0.5 as data augmentations. As depicted in Figure 7, we observe that compared with mixup, random crop and random flip result in a more evident vertical shift in the performance curve.